OpenStudy (anonymous):

Find all solutions 7 sin2x - 14 sin x + 2 = -5

OpenStudy (tkhunny):

Have you considered adding 5 to both sides and then: 1) Factoring, if you can: \((7\sin(x) - What)(\sin(x) - What)\) 2) Quadratic formula: \(\sin(x) = \dfrac{14\pm\sqrt{Stuff}}{2\cdot 7}\) Ringing any bells?

OpenStudy (anonymous):

Simplify your equation, \[7\sin^2x-14sinx+7=0\] Replace sinx by say, t \[7t^2-14t+7=0\] Remove a factor of 7,\[t^2-2t+1=0\] Solve using quadratic equation and in the end replace your t by sin(x) again Then take inverse sine on both sides to solve for x